OpenStudy (anonymous):

Freida drove 20 miles in 24 minutes. At this rate, how many miles did she drive in 6 minutes?

OpenStudy (anonymous):

@poual-39816

OpenStudy (twolsey5):

5 miles

jabez177 (jabez177):

4 maybe but your friend can check

OpenStudy (bookworm14):

so you are looking for how far she can go in 6 minutes vs 24 minutes....in order to get to 6 from 24 you must divide 24 by 4... 24/4=6...so you have to do the same thing to the other side...20/4=5.... she drve 5 miles in 6 minutes...Hope this helps!! ~Bree :)

OpenStudy (twolsey5):

4*6=20 4 miles times 6 minutes = 20 miles

jabez177 (jabez177):

wow smart that also helped me

OpenStudy (anonymous):

so its 5 miles?

OpenStudy (twolsey5):

hey can some one help me with my question pls this is the link http://openstudy.com/study#/updates/547ca862e4b082a9810fa6e3

OpenStudy (bookworm14):

yes the answer is 5 miles

OpenStudy (anonymous):

ok

OpenStudy (bookworm14):

@twolsey5 i think you made an error in your math you put "4*6=20" when 4*6 actually equals 24 (i know that is what you meant to put, you just typed it wrong xD )

OpenStudy (anonymous):

ok